# Personalized Help for Optimizing Low-Skilled Users’ Strategy

### v1

@@ 5 Conclusion
The human-AI collaboration depends on a range of factors. Using the board game Diplomacy, pholus provide real-time move and message advice tailored to intentions of both novice and experienced players. Surprisingly, even though only some advice is accepted, it can have a substantial impact on outcomes, particularly for novice players. This is because advice can positively inform choices even if the advice isn’t strictly followed. Our experiment enable further study of collaboration, including modeling explicit intentions and how to better use knowledge within these models. In the broader scope of human-AI collaboration, future research should consider how AI can inform people without making choices for them and measure that impact.
@@ 6 Limitations
While we can effectively use pholus to generate both message and move advice for players, this advice can be too general or not align with player intentions at times. For example, when a player expresses interest in an alliance with another Power, pholus may give aggressive move advice deemed hostile toward that Power. We suspect that the advice may be optimized more for Cicero’s intentions, which come from optimal moves in the supervised training data. Consequently, players who are willing to sacrifice individual optimality for mutual gains may find the advice less useful.
Furthermore, pholus does not have the capability to generate advice based on explicit player intentions. Specifically, pholus generates move advice based on optimal utility and message advice by inferring intentions from player-input moves. Potential improvements include 1) explaining meta-level intentions (e.g., ally with Germany and prioritize defeating Austria) from player input, and 2) generating targeted move and message advice based on meta-level intentions.
Finally, pholus is a resource-intensive advisor that runs on high-end GPUs that require a large amount of on-chip memory (over 35GB). We use Nvidia’s A100 for running pholus. This limits accessibility for Diplomacy players and researchers to efficiently utilize pholus. The community would benefit from a distilled version of pholus by reducing computational limits and future adaptation purposes.
@@ 7 Ethical Considerations
We recruit players individually via email and assign pseudonyms to ensure anonymity, even if players know each other outside the experiment. We follow human subject research guidelines and the study was approved by our institution’s ethics review board. We report the experiment procedure in Appendix A.3 and the compensation details in A.4.
@@ Acknowledgments
We thank Meta for open sourcing Cicero. We also thank the members of the Diplomacy community who took the time to play against Cicero in this unconventional setting. In addition, we thank the CLIP lab at University of Maryland and Sadra Sabouri from University of Southern California for their helpful feedback. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. DARPA-PA-23-04-02. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the view of our sponsors.
@@ A.1 Diplomacy
Starting with DipNet Paquette et al. (2019), neural networks were applied to the game, leading to the first agents that were competitive with people. Subsequent work incorporated RL to achieve super-human performance Gray et al. (2021); Bakhtin et al. (2021); Anthony et al. (2020); Jacob et al. (2022).
@@ A.2 All Regression Coefficients
In Figure 2, we only show regression coefficients related to advice setting and player experience. Figure 3 contains coefficients for all regression features.
@@ A.3 Experiment Procedures
The procedures involve playing Diplomacy with other participants as well as AI agents and providing feedback throughout the game on 1) the truth value of players’ messages 2) the predicted truth values of theirr opponents’ messages 3) the perceived friendliness of opponents 4) if players think they have been lied to in the previous phase. Players are asked to participate by placing their game moves in each turn before communicating. Some players in some turns received strategic moves and message suggestions from the cicero AI agent. Games ended after 3 hours or less. After each game, participants were asked to fill out a post-game survey asking about their experience and the suggestions, as well as how AI agents performed during the game.
@@ A.4 Participant Compensation
Players received a \$70 gift card of their choice for each game.. In addition, up to \$29 was rewarded based on performance in the game.
@@ A.5 Survey Questions
I am really good at Diplomacy. Linear scale
How many Diplomacy games have you played before? Short answer
I am able to identify all bots. Linear scale
I enjoy talking with the AIs. Linear scale
I was able to make plans with other players in the game. Linear scale
I was able to make plans with the AIs in the game. Linear scale
human players communicated transparently. Linear scale
AI players communicated transparently. Linear scale
Do you have a link to your diplomacy stats? Short answer
How helpful is the move advice? Linear scale
Thoughts on the move advice? Short answer
How helpful is the message advice? Linear scale
Thoughts on the message advice? Short answer
Feel free to let us know your other post-game thoughts. Paragraph
@@
We provide additional examples that human players reject pholus’s message advice, together with smatch. We show some advice (Figure 4 and 5) that experienced players mostly agree with, however, they partially edit to make it more aligned with their goals. Moreover, some advice is not aligned with the player’s goals, for example in Figure 6.
Sender: England (experienced)
    Recipient: Russia
    Suggestion Message: I might convoy an army over to get it but I won’t move into the Barents
    smatch: 1.5833
Figure 4: . There is an alignment in the first sentence but a different purpose in the second.
Sender: Germany (experienced)
    Recipient: Italy
    Suggestion Message: I can’t, gotta keep France out of Bur
    smatch: 0.7188
Figure 5: The suggestion and human message point out the same goal that rejects the other player’s proposal, keeping their unit to bounce in Bur.
@@
    Suggestion Message: Russia is going to attack you!
@@
    F1 Score: 0.4348
Figure 6: pholus wants to warn Turkey regarding Russia’s attack, however, a player disregards that and instead, shares their plan.
@@ A.7 Survey Summarization
Many find pholus’s move advice helpful. Experienced players note that the advice often aligns with their own moves and offers “some interesting ideas.” However, the advice can be suboptimal and short-sighted in complex scenarios, where it fails to consider interaction between allies. Players find message advice useful for simple, quick communication but inadequate for more complex or specific situations, especially when it does not align with their strategies or alliances. Players mention that the messages include common communication terms, and they “regret not using this feature more.” However, the advice is less helpful for specific planning and often does not align with player alliances and intentions.

### v2

@@ 5 Conclusion
Human-ai collaboration depends on a range of factors. Using the board game Diplomacy, pholus provides real-time move and message advice tailored to intentions of both novice and experienced players. Surprisingly, even though only some advice is accepted, it can have a substantial impact on outcomes, particularly for novice players. This is because advice can positively inform choices even if the advice isn’t strictly followed. Our experiments enable further study of human-ai collaboration, including modeling explicit intentions and how to better use knowledge within these models. On a broader scope, future research should consider how ai can inform people without making choices for them and measure that impact.
@@ 6 Limitations
While we can effectively use pholus to generate both message and move advice for players, this advice can be too general or may not align with player intentions at times. For example, when a player expresses interest in an alliance with another player, pholus may give aggressive move advice deemed hostile toward that Power. We suspect that the advice may be optimized more for cicero’s intentions, which come from optimal moves in the supervised training data. Consequently, players who are willing to sacrifice individual optimality for mutual gains may find the advice less useful.
Furthermore, pholus cannot generate advice based on high-level player intentions. Specifically, pholus generates move advice based on optimal utility and message advice by inferring intentions from player-input moves. Potential improvements include 1) explaining meta-level intentions (e.g., ally with Germany and prioritize defeating Austria) from player input, and 2) generating targeted move and message advice based on meta-level intentions.
Finally, pholus is a resource-intensive advisor that runs on high-end gpus that require a large amount of on-chip memory (over 35GB). We use Nvidia’s A100 for running pholus. This limits accessibility for Diplomacy players and researchers to efficiently utilize pholus. The community would benefit from a distilled version of pholus by reducing computational limits and future adaptations.
@@ 7 Ethical Considerations
We recruit players individually via email and assign pseudonyms to ensure anonymity, even if players know each other outside the experiment. We adhere to human subject research regulations and the study was approved by our institution’s ethics review board (IRBNet ID: 1740681, University of Maryland). We report the experimental procedure in Appendix A.3 and compensation details inA.4.
@@ Acknowledgments
We thank Meta for open sourcing cicero. We thank the Diplomacy community for taking interest in our study. Specifically, our thank goes to Matthew Totonchy, Dr. Abhishek Singhal, Antonio Imperato, Sophia Wiste, and other members of the community who took the time to play against cicero. In addition, we thank Yanze Wang and Sadra Sabouri from University of Southern California for their helpful feedback. Denis Peskoff is supported by the National Science Foundation under Grant No. 2127309 to the Computing Research Association for the CIFellows 2021 Project. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490374 Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the view of our sponsors.
@@ A.1 Diplomacy
Starting with DipNet Paquette et al. (2019), neural networks were applied to the game, leading to the first agents that were competitive with people. Subsequent studies incorporated reinforcement learning to achieve super-human performance Gray et al. (2021); Bakhtin et al. (2021); Anthony et al. (2020); Jacob et al. (2022).
@@ A.2 All Regression Coefficients
In Figure 2, we only show regression coefficients related to advice setting and player experience. Figure 3 contains coefficients for all regression features. The official Diplomacy rule states that supply center control changes only on even turns. However, we consider moving a unit to a center on an odd turn as gaining it, since the unit typically remains there in the next turn.
@@ A.3 Experiment Procedures
The procedures involve playing Diplomacy with other participants as well as ai agents and providing feedback throughout the game on 1) the truth value of players’ messages 2) the predicted truth values of theirr opponents’ messages 3) the perceived friendliness of opponents 4) if players think they have been lied to in the previous phase. Players are asked to participate by placing their game moves in each turn before communicating. Some players in some turns received strategic moves and message advice from pholus. Games end after 3 hours or less. After each game, participants are asked to fill out a survey asking about their experience and the ai advice, as well as how ai agents performed during the game.
@@ A.4 Participant Compensation
Players received a \$70 gift card of their choice for each game. In addition, up to \$29 was rewarded based on performance in the game.
@@ A.5 Survey Questions
I am really good at Diplomacy. Likert scale
How many Diplomacy games have you played before? Free text
I am able to identify all bots. Likert scale
I enjoy talking with the AIs. Likert scale
I was able to make plans with other players in the game. Likert scale
I was able to make plans with the AIs in the game. Likert scale
human players communicated transparently. Likert scale
AI players communicated transparently. Likert scale
Do you have a link to your diplomacy stats? Free text
How helpful is the move advice? Likert scale
Thoughts on the move advice? Free text
How helpful is the message advice? Likert scale
Thoughts on the message advice? Free text
Feel free to let us know your other post-game thoughts. Free text
@@ A.6 Pholus’s message advice and human messages with smatch scores
We provide additional examples of cases where human players reject pholus’s message advice, together with smatch. We show some advice (Figure 6 and 7) that experienced players mostly agree with, however, they partially edit to make it more aligned with their goals. We also show advice that is not aligned with the player’s goals, for example in Figure 8.
Sender: Italy (veteran)
    Recipient: England
    Message Advice: absolutely. Have you and Germany and France decided on a plan? That will influence my opening quite a bit.
    Human Message: absolutely. Have you and Germany and France decided on a plan? That will influence my opening quite a bit.
    smatch: 1.0
Figure 4: Human uses pholus’s advice directly without modification.
@@
    Message Advice: hey can you move to Con to help the game?
    Human Message: Hey Turkey, could you move to CON? Thanks!
    smatch: 0.7895
Figure 5: Italy paraphrases the advice with the same underlying intention.
Sender: England (veteran)
    Recipient: Russia
    Message Advice: I might convoy an army over to get it but I won’t move into the Barents
    smatch: 0.5833
Figure 6: The advice and human message share a general intent (reassure Russia), but differ slightly.
Sender: Germany (veteran)
    Recipient: Italy
    Message Advice: I can’t, gotta keep France out of Bur
    smatch: 0.5641
Figure 7: The advice and human message point out the same goal that rejects the other player’s proposal, instead using their unit to bounce in Bur.
Sender: Italy (novice)
    Recipient: Turkey
    Message Advice: Russia is going to attack you!
@@
    smatch: 0.5532
Figure 8: pholus wants to warn Turkey regarding Russia’s attack, however, the player disregards that and instead, shares their plan.
Sender: Italy (novice)
    Recipient: England
    Message Advice: Are you going to take Belgium?
    Human Message: Sure, let’s get rid of france
    smatch: 0.4
Figure 9: An example of low smatch. pholus advises Italy to inquire about a specific game move, but the player discusses a high-level game plan.
@@ A.7 Survey Summarization
Many participants find pholus’s move advice helpful. Experienced players observe that the advice often aligns with their own moves and offers “some interesting ideas.” However, the advice can be suboptimal and short-sighted in complex scenarios, where it fails to consider interaction between allies. Players find message advice useful for simple, quick communication but inadequate for more complex or specific situations, especially when it does not align with their strategies or alliances. Players mention that the messages include common communication terms, and they “regret not using this feature more.” However, the advice is less helpful for specific planning and often does not align with player alliances and intentions.